# The Taylor-Wiles method for coherent cohomology, II

### v2

@@ Introduction
The reader should understand that, in contrast to most papers on the Taylor-Wiles method, our aim is not to show that certain Galois representations are automorphic (or potentially automorphic). This is already known for the kinds of Galois representations that arise in the cohomology of the PEL Shimura varieties that are the subject of these papers. In contrast, the primary purpose of the present paper, as well as its predecessor [27], is to obtain results about the structure of the localized Hecke algebra ${\mathbb{T}}$, with a view to applications to arithmetic problems. Thus the main results of the paper are those contained in §4 and §5. Thus, let $F$ be a CM field in which every prime dividing $p$ splits over the maximal totally real subfield $F^{+}$ of $F$, and let $\pi$ be a cuspidal cohomological automorphic representation of $GL(n)_{F}$. When $V$ is an $n$-dimensional hermitian space over $F$, and $G$ its unitary similitude group, we consider the $L$-packet $\Pi_{V}(\pi)$ of automorphic representations of discrete series type that base change to $\pi$; we temporarily denote by $S_{V}$ the corresponding Shimura variety. We specifically prove – under the hypotheses that allow us to apply the Taylor-Wiles method – that
@@ 1.1.1. Type data
It was pointed out in [27, Remark 6.9] that the hypotheses of that paper restricted ramification of the residual representation $\bar{\rho}$ at places of $F^{+}$ not dividing $p$. In this paper those restrictions are somewhat relaxed, although we only consider deformation problems that involve minimal ramification. Let $S=S(K)$ be the set of primes $q$ where either $K_{v}$, for some $v\in D_{+}(q)$, or $K_{-,q}$, is not hyperspecial maximal; we call $S$ the set of ramified primes for $K$. The set $S$ contains the set $S(G)$ of primes $q$ at which the group $G$ is ramified, and in particular contains the rational primes divisible by (the set) $S(F^{+})$ of primes of $F^{+}$ that ramify in $F/F^{+}$. At primes $q\in S(G)$ we take $K_{-,q}$ to be special maximal. If $q\in S\setminus S(G)$ we assume $K_{-,q}$ is a hyperspecial maximal subgroup of $G_{-,q}$. For $v\in D_{+}(q)$, whether or not $q\in S(G)$, we will choose a set of type data, namely quadruples $(K_{v}^{+},K_{v},\mbox{ю}_{v},\Lambda_{\mbox{ю}_{v}})$, with $K_{v}\subset K^{+}_{v}$ a pair of open compact subgroups, $\Lambda_{\mbox{ю}_{v}}$ a finite free ${\mathcal{O}}$-module and
@@
where ${\mathfrak{m}}_{{\mathfrak{r}}_{0}}$ is the maximal ideal in the integer ring and $u$ is upper triangular unipotent. This suffices to guarantee that $K$ is neat, and ${\mathfrak{r}}_{0}$ will be chosen so that the deformation problems to be studied below are minimal and unrestricted at ${\mathfrak{r}}$. In the statement of [27, Theorem 6.8] this set is called $S_{a}$. [^2] Erratum for [27]: In the published version of [27] the set $S_{a}$ was mentioned appropriately in the statement of its main theorem but its definition was inadvertently omitted. As already mentioned, this condition is required in order to guarantee smoothness. Theorem 36 requires as well that the deformation condition at such an ${\mathfrak{r}}$ be unrestricted and minimal. We choose ${\mathfrak{r}}$ as on [10, p. 111] (where it is called $S_{a}$) or as on [60, p. 893] (where it is called $v_{1}$). The hypothesis on ${\mathfrak{r}}$ – that $[F(\zeta_{\mathfrak{r}}):F]>n$ – guarantees that ${\mathfrak{r}}$ satisfies condition (1) of Theorem 36.
@@ 1.1.2. Notation for highest weights
Notation is as in [27, §2]. We choose a maximal torus $T\subset K_{x}$; then $T$ is also a maximal torus in $G$. Let $\Phi=\Phi(G,T)$ denote the set of roots of $G$ relative to $T$, and let $\Phi^{+}$ be a system of positive roots compatible with $P_{x}$, i.e. containing the roots of $\mathfrak{p}_{x}^{-}$; let $X^{+}(T)$ (resp. $X_{x}^{+}(T)$) denote the set of dominant weights for $G$ (resp. for $K_{x}$) relative to this choice. Let $W=W(G,T)$ be the absolute Weyl group and let $W^{x}\subset W$ be the set of Kostant representatives.
@@
(cf. [47, 2.32]) Let $\mu\in X^{+}(T)$. Say $\mu$ is $p$-small, resp. $p$-small relative to $x$, if
@@
in the notation of [47, Definition 3.2]. (In [47] this is called “$p$-small for the geometric realization of Weyl’s construction.")
@@ 1.3. Terminology for toroidal compactifications
(Lan) Let $p$ be a prime at which $G$ is unramified, and fix a hyperspecial maximal compact $K_{p}\subseteq G({\mathbb{Q}}_{p}).$ Then for all neat $K\supset K_{p}$, there is a compatible choice of admissible smooth rational polyhedral data $\Sigma$ (see [42, Def. 6.3.3.2]) such that the ${\mathbb{S}}_{K}(G,X)_{\Sigma}$ is a smooth proper scheme over a $p$-adic integer ring ${\mathcal{O}}$ which contains ${\mathbb{S}}_{K}(G,X)$ of Proposition 10 as an open dense subscheme. Furthermore, $\Sigma$ may be chosen so that ${\mathbb{S}}_{K}(G,X)_{\Sigma}-{\mathbb{S}}_{K}(G,X)$, viewed as a closed reduced subscheme, is a divisor with normal crossings. There is a canonical strata-preserving isomorphism between the basechange to ${\mathrm{Spec}}({\mathbb{C}})$ of this integral model and the classical analytical construction as in [2].
@@ 1.3.1. Parabolic strata
The boundary divisor $\partial S_{K}(G,X)_{\Sigma}$ has a closed covering indexed by maximal standard rational parabolic subgroups of $G$, defined as follows. Let ${\mathbb{S}}_{K}^{min}$ denote the minimal compactification of ${\mathbb{S}}_{K}$ over $Spec({\mathcal{O}})$, as in [42, Theorem 7.4.2.1]. The boundary ${\mathbb{S}}_{K}^{min}\setminus{\mathbb{S}}_{K}$ decomposes as a disjoint union of locally closed strata $\partial^{P}{\mathbb{S}}_{K}^{min}$ indexed by standard maximal rational parabolic subgroups $P\subset G$. Each $\partial^{P}{\mathbb{S}}_{K}^{min}$ is a union of certain strata denoted $\mathbf{Z}_{[(\Phi_{\mathcal{H}},\delta_{\mathcal{H}})]}$ in Lan, where $P$ is the stabilizer of the filtration on the hermitian vector space $V$ concealed in the notation $(\Phi_{\mathcal{H}},\delta_{\mathcal{H}})$ (see [42, §5.4] for an explanation).
There is a canonical morphism [42, Theorem 7.4.2.1, 3] $\oint:{\mathbb{S}}_{K,\Sigma}{\leavevmode\nobreak\ \rightarrow\leavevmode\nobreak\ }{\mathbb{S}}_{K}^{min}$. For each standard maximal rational parabolic $P\subset G$, we let the $P$-stratum $\partial^{P}{\mathbb{S}}_{K,\Sigma}$ of the toroidal boundary of ${\mathbb{S}}_{K,\Sigma}$ be the closure of $\oint^{-1}(\partial^{P}{\mathbb{S}}_{K}^{min})$. Let $R\subset G$ be a standard rational parabolic, and write $R=\cap Q_{j}$ for a (unique) set of standard maximal parabolics $Q_{j}$. We define the $R$-stratum:
@@
Suppose $R$ has parabolic rank $r$. We define the nerve $\mathfrak{N}_{\Sigma}(R)$ of the closed covering of $\partial^{R}{\mathbb{S}}_{K,\Sigma}$ by irreducible components of codimension $r$, as in [33, §3.1] (where the $R$-stratum was denoted $Z_{\Sigma}(R)$). We recall the relation between the homotopy type of $\mathfrak{N}_{\Sigma}(R)$ and Borel-Serre compactifications in §8.2.1.
@@ 1.4. Level subgroups
The following theorem is proved in §A.4 of the first author’s thesis.
@@
Suppose $K=\prod_{q}K_{q}\subseteq G({{\mathbf{A}}_{f}})$ is a neat subgroup, such that $K_{q}$ are maximal hyperspecial for almost all $q$. Suppose further that $K\supset K_{p}$, a fixed hyperspecial maximal $K_{p}\subseteq G({\mathbb{Q}}_{p}).$ Fix a set $Q$ of primes as above. Let ${\mathcal{B}}$ be the set of modifications $K_{i,Q}(t)$ of $K_{i,Q}$, with $i=0,1,\Delta$. Then for every $K_{i,Q}(t)\in{\mathcal{B}}$ there is a toroidal datum $\Sigma(K_{i,Q}(t))$ such that
@@
    If $K^{\prime}\subset K$ is an inclusion in ${\mathcal{B}}$ then the natural covering map $S_{K^{\prime}}(G,X){\leavevmode\nobreak\ \rightarrow\leavevmode\nobreak\ }S_{K}(G,X)$ extends to a map
@@
    For any $w$ and any $t\in G_{w}$ as above, the map
    $S_{K_{\Delta,Q}(t)}(G,X)_{\Sigma_{\Delta,Q}(t)}{\leavevmode\nobreak\ \rightarrow\leavevmode\nobreak\ }S_{K_{0,Q}(t)}(G,X)_{\Sigma_{0,Q}(t)}$
@@
    Conditions (1)-(3) hold when the $S_{K_{i,Q}(t)}(G,X)$ are replaced by their integral models ${\mathbb{S}}_{K_{i,Q}(t)}$.
The first two assertions follow from the classical theory of toroidal compactifications. The third one requires more delicate analysis of the analytic construction in [2] and uses the observation that $U_{F}({\mathbb{Q}}_{q})$ is contained in a pro-$q$ group for each boundary component $F$ and unramified prime $q,$ while $\Delta_{Q}$ is a $p$-group to show that $K_{\Delta,Q}(t)$ and $K_{0,Q}(t)$ restrict to the same subgroup of $U_{F}({\mathbb{Q}}).$ The last claim follows by purity of the branch locus applied to each irreducible boundary component which reduces it to the characteristic zero statement in (3).
@@ 2.1. Adelic representations and fixed vectors
(21) $E_{2}^{r,s}=H^{r}(K,R^{s}H({\mathcal{F}}_{K_{p}}^{?})\Rightarrow H^{r+s}({\mathbb{S}}_{K,\Sigma_{K}},{\mathcal{F}}_{K}^{?}),$
@@
[6] With notation as above, $H^{r}(K,R^{s}H({\mathcal{F}}_{K_{p}}^{?})=0$ for all $r>0$, ?$=can,sub$. Thus the canonical map
$RH({\mathcal{F}}_{K}^{?})\rightarrow RHom_{K^{p}}(1,RH({\mathcal{F}}_{K_{p}}^{?}),$
@@ 2.2. Hecke operators
Notation is as in Proposition 17. Fix $K$ hyperspecial at $p$ and let $S=S(K)$ be as above. For $t\in G({{\mathbf{A}}_{f}^{S}})$, a Hecke operator $T(t)\in{\mathrm{End}}(RH({\mathcal{F}}_{K}^{?}))$, denoted $T_{t}^{\Sigma}$, is defined in [17, §8.1.5], for ?$=can,sub$. An alternative definition is given in [7, §4.2], where it is also proved that the subalgebra $T_{S,K}$ of ${\mathrm{End}}(RH({\mathcal{F}}_{K}^{?}))$ generated by the $T(t)$ is commutative. The algebra $T_{S,K}$ naturally maps to ${\mathrm{End}}(\oplus_{i}H^{i}(RH({\mathcal{F}}_{K})))$ but the map to its image is not necessarily an isomorphism. In particular, there is no reason to assume $T_{S,K}$ to be reduced.
@@
Let $\nu$ be as in Theorem 24, with $k$ a field of characteristic $p$. Let ${\mathfrak{m}}=\ker{\nu}$; we write ${\mathfrak{m}}={\mathfrak{m}}_{\nu}$ or ${\mathfrak{m}}={\mathfrak{m}}_{\rho}$ with $\rho=\rho_{\nu}$. Let ${\mathbb{T}}_{\nu}$ be the reduced localization of $T_{S,K}$ at ${\mathfrak{m}}_{\nu}$, and let
@@
    $\mu$ is sufficiently regular in the sense of [47, 7.18]
@@
The Lan-Suh vanishing conditions are written out more explicitly in [27, §6.10]. We write $H^{q}_{!}({\mathbb{S}},{\mathcal{F}}_{K})$ instead of $H^{q}_{!}(R\Gamma({\mathbb{S}},{\mathcal{F}}_{K}))$.

### v3

@@ Introduction
The reader should understand that, in contrast to most papers on the Taylor-Wiles method, our aim is not to show that certain Galois representations are automorphic (or potentially automorphic). This is already known for the kinds of Galois representations that arise in the cohomology of the PEL Shimura varieties that are the subject of these papers. In contrast, the primary purpose of the present paper, as well as its predecessor [28], is to obtain results about the structure of the localized Hecke algebra ${\mathbb{T}}$, with a view to applications to arithmetic problems. Thus the main results of the paper are those contained in §4 and §5. Thus, let $F$ be a CM field in which every prime dividing $p$ splits over the maximal totally real subfield $F^{+}$ of $F$, and let $\pi$ be a cuspidal cohomological automorphic representation of $GL(n)_{F}$. We assume the contragredient of $\pi$ is isomorphic to $\pi^{c}$, where $c$ is the non-trivial element of $Gal(F/F^{+})$. When $V$ is an $n$-dimensional hermitian space over $F$, and $G$ its unitary similitude group, we consider the $L$-packet $\Pi_{V}(\pi)$ of automorphic representations of discrete series type that base change to $\pi$, and we assume $\Pi_{V}(\pi)$ is non-empty; the conjugate duality hypothesis is a necessary condition for this, but there may also be local conditions at places where $G$ is not quasi-split (see [42]). We temporarily denote by $S_{V}$ the corresponding Shimura variety. We specifically prove – under the hypotheses that allow us to apply the Taylor-Wiles method – that
@@ 1.1.1. Type data
It was pointed out in [28, Remark 6.9] that the hypotheses of that paper restricted ramification of the residual representation $\bar{\rho}$ at places of $F^{+}$ not dividing $p$. In this paper those restrictions are somewhat relaxed, although we only consider deformation problems that involve minimal ramification. Let $S=S(K)$ be the set of primes $q$ where either $K_{v}$, for some $v\in D_{+}(q)$, or $K_{-,q}$, is not hyperspecial maximal; we call $S$ the set of ramified primes for $K$. The set $S$ contains the set $S(G)$ of primes $q$ at which the group $G$ is ramified, and in particular contains the rational primes divisible by (the set) $S(F^{+})$ of primes of $F^{+}$ that ramify in $F/F^{+}$. At primes $q\in S(G)$ we take $K_{-,q}$ to be special maximal. If $q\in S\setminus S(G)$ we assume $K_{-,q}$ is a hyperspecial maximal subgroup of $G_{-,q}$. For $v\in D_{+}(q)$, whether or not $q\in S(G)$, we will choose a set of type data, namely quadruples $(K_{v}^{+},K_{v},\mbox{ю}_{v},\Lambda_{\mbox{ю}_{v}})$, with $K_{v}\subset K^{+}_{v}$ a pair of open compact subgroups, $\Lambda_{\mbox{ю}_{v}}$ a finite free ${\mathcal{O}}$-module and
@@
where ${\mathfrak{m}}_{{\mathfrak{r}}_{0}}$ is the maximal ideal in the integer ring and $u$ is upper triangular unipotent. This suffices to guarantee that $K$ is neat, and ${\mathfrak{r}}_{0}$ will be chosen so that the deformation problems to be studied below are minimal and unrestricted at ${\mathfrak{r}}$. In the statement of [28, Theorem 6.8] this set is called $S_{a}$. [^2] Erratum for [28]: In the published version of [28] the set $S_{a}$ was mentioned appropriately in the statement of its main theorem but its definition was inadvertently omitted. As already mentioned, this condition is required in order to guarantee smoothness. Theorem 36 requires as well that the deformation condition at such an ${\mathfrak{r}}$ be unrestricted and minimal. We choose ${\mathfrak{r}}$ as on [11, p. 111] (where it is called $S_{a}$) or as on [62, p. 893] (where it is called $v_{1}$). The hypothesis on ${\mathfrak{r}}$ – that $[F(\zeta_{\mathfrak{r}}):F]>n$ – guarantees that ${\mathfrak{r}}$ satisfies condition (1) of Theorem 36.
@@ 1.1.2. Notation for highest weights
Notation is as in [28, §2]. We choose a maximal torus $T\subset K_{x}$; then $T$ is also a maximal torus in $G$. Let $\Phi=\Phi(G,T)$ denote the set of roots of $G$ relative to $T$, and let $\Phi^{+}$ be a system of positive roots compatible with $P_{x}$, i.e. containing the roots of $\mathfrak{p}_{x}^{-}$; let $X^{+}(T)$ (resp. $X_{x}^{+}(T)$) denote the set of dominant weights for $G$ (resp. for $K_{x}$) relative to this choice. Let $W=W(G,T)$ be the absolute Weyl group and let $W^{x}\subset W$ be the set of Kostant representatives.
@@
(cf. [49, 2.32]) Let $\mu\in X^{+}(T)$. Say $\mu$ is $p$-small, resp. $p$-small relative to $x$, if
@@
in the notation of [49, Definition 3.2]. (In [49] this is called “$p$-small for the geometric realization of Weyl’s construction.")
@@ 1.3. Terminology for toroidal compactifications
(Lan) Let $p$ be a prime at which $G$ is unramified, and fix a hyperspecial maximal compact $K_{p}\subseteq G({\mathbb{Q}}_{p}).$ Then for all neat $K\supset K_{p}$, there is a compatible choice of admissible smooth rational polyhedral data $\Sigma$ (see [44, Def. 6.3.3.2]) such that the ${\mathbb{S}}_{K}(G,X)_{\Sigma}$ is a smooth proper scheme over a $p$-adic integer ring ${\mathcal{O}}$ which contains ${\mathbb{S}}_{K}(G,X)$ of Proposition 10 as an open dense subscheme. Furthermore, $\Sigma$ may be chosen so that ${\mathbb{S}}_{K}(G,X)_{\Sigma}-{\mathbb{S}}_{K}(G,X)$, viewed as a closed reduced subscheme, is a divisor with normal crossings. There is a canonical strata-preserving isomorphism between the basechange to ${\mathrm{Spec}}({\mathbb{C}})$ of this integral model and the classical analytical construction as in [2].
@@ 1.3.1. Parabolic strata
The boundary divisor $\partial S_{K}(G,X)_{\Sigma}$ has a closed covering indexed by maximal standard rational parabolic subgroups of $G$, defined as follows. Let ${\mathbb{S}}_{K}^{min}$ denote the minimal compactification of ${\mathbb{S}}_{K}$ over $Spec({\mathcal{O}})$, as in [44, Theorem 7.4.2.1]. The boundary ${\mathbb{S}}_{K}^{min}\setminus{\mathbb{S}}_{K}$ decomposes as a disjoint union of locally closed strata $\partial^{P}{\mathbb{S}}_{K}^{min}$ indexed by standard maximal rational parabolic subgroups $P\subset G$. Each $\partial^{P}{\mathbb{S}}_{K}^{min}$ is a union of certain strata denoted $\mathbf{Z}_{[(\Phi_{\mathcal{H}},\delta_{\mathcal{H}})]}$ in Lan, where $P$ is the stabilizer of the filtration on the hermitian vector space $V$ concealed in the notation $(\Phi_{\mathcal{H}},\delta_{\mathcal{H}})$ (see [44, §5.4] for an explanation).
There is a canonical morphism [44, Theorem 7.4.2.1, 3] $\oint:{\mathbb{S}}_{K,\Sigma}{\leavevmode\nobreak\ \rightarrow\leavevmode\nobreak\ }{\mathbb{S}}_{K}^{min}$. For each standard maximal rational parabolic $P\subset G$, we let the $P$-stratum $\partial^{P}{\mathbb{S}}_{K,\Sigma}$ of the toroidal boundary of ${\mathbb{S}}_{K,\Sigma}$ be the closure of $\oint^{-1}(\partial^{P}{\mathbb{S}}_{K}^{min})$. Let $R\subset G$ be a standard rational parabolic, and write $R=\cap Q_{j}$ for a (unique) set of standard maximal parabolics $Q_{j}$. We define the $R$-stratum:
@@
Suppose $R$ has parabolic rank $r$. We define the nerve $\mathfrak{N}_{\Sigma}(R)$ of the closed covering of $\partial^{R}{\mathbb{S}}_{K,\Sigma}$ by irreducible components of codimension $r$, as in [34, §3.1] (where the $R$-stratum was denoted $Z_{\Sigma}(R)$). We recall the relation between the homotopy type of $\mathfrak{N}_{\Sigma}(R)$ and Borel-Serre compactifications in §8.2.1.
@@ 1.4. Level subgroups
The following theorem is proved in §A.6 of the first author’s thesis [3].
@@
Suppose $K=\prod_{q}K_{q}\subseteq G({{\mathbf{A}}_{f}})$ is a neat subgroup, such that $K_{q}$ are maximal hyperspecial for almost all $q$. Suppose further that $K\supset K_{p}$, a fixed hyperspecial maximal $K_{p}\subseteq G({\mathbb{Q}}_{p}).$ Fix a set $Q$ of primes as above. Let ${\mathcal{B}}$ be the set of modifications $K_{i,Q}(t)$ of $K_{i,Q}$, with $i=0,1,\Delta$. Then for every $K_{i,Q}(t)$ there is a toroidal datum $\Sigma(K_{i,Q}(t))$ such that
@@
    If $K^{\prime}\subset K$ is an inclusion in ${\mathcal{B}}$ then the natural covering map
    $S_{K^{\prime}}(G,X){\leavevmode\nobreak\ \rightarrow\leavevmode\nobreak\ }S_{K}(G,X)$
    extends to a map
@@
    The map
    $S_{K_{\Delta,Q}}(G,X)_{\Sigma_{\Delta,Q}}{\leavevmode\nobreak\ \rightarrow\leavevmode\nobreak\ }S_{K_{0,Q}}(G,X)_{\Sigma_{0,Q}}$
@@
    Conditions (1)-(3) hold when the $S_{K_{i,Q}}(G,X)$ are replaced by their integral models ${\mathbb{S}}_{K_{i,Q}}$.
The first two assertions follow from the classical theory of toroidal compactifications and are used without comment in defining Hecke correspondences in §2.2. The third one requires more delicate analysis of the analytic construction in [2]. Atanasov uses the observation that the intersection of $U_{F}({\mathbb{Q}}_{q})$ with any compact open subgroup is contained in a pro-$q$ group for each $F$ and each unramified prime $q\neq p$. Since $\Delta_{Q}$ is a $p$-group, it then suffices to show that $K_{\Delta,Q}$ and $K_{0,Q}$ restrict to the same subgroup of $U_{F}({\mathbb{Q}})$, and this is a simple calculation. The last claim follows by purity of the branch locus applied to each irreducible boundary component, which reduces it to the characteristic zero statement in (3).
@@ 2.1. Adelic representations and fixed vectors
(21) $E_{2}^{r,s}=H^{r}(K,R^{s}H({\mathcal{F}}_{K_{p}}^{?}))\Rightarrow H^{r+s}({\mathbb{S}}_{K,\Sigma_{K}},{\mathcal{F}}_{K}^{?}),$
@@
[7] With notation as above, $H^{r}(K,R^{s}H({\mathcal{F}}_{K_{p}}^{?}))=0$ for all $r>0$, ?$=can,sub$. Thus the canonical map
$RH({\mathcal{F}}_{K}^{?})\rightarrow RHom_{K^{p}}(1,RH({\mathcal{F}}_{K_{p}}^{?})),$
@@ 2.2. Hecke operators
Notation is as in Proposition 17. Fix $K$ hyperspecial at $p$ and let $S=S(K)$ be as above. For $t\in G({{\mathbf{A}}_{f}^{S}})$, a Hecke operator $T(t)\in{\mathrm{End}}(RH({\mathcal{F}}_{K}^{?}))$, denoted $T_{t}^{\Sigma}$, is defined in [18, §8.1.5], for ?$=can,sub$. An alternative definition is given in [8, §4.2], where it is also proved that the subalgebra $T_{S,K}$ of ${\mathrm{End}}(RH({\mathcal{F}}_{K}^{?}))$ generated by the $T(t)$ is commutative. The algebra $T_{S,K}$ naturally maps to ${\mathrm{End}}(\oplus_{i}H^{i}(RH({\mathcal{F}}_{K})))$ but the map to its image is not necessarily an isomorphism. In particular, there is no reason to assume $T_{S,K}$ to be reduced.
@@
Let $\nu$ be as in Theorem 24, with $k$ a field of characteristic $p$. Let ${\mathfrak{m}}=\ker{\nu}$; we write ${\mathfrak{m}}={\mathfrak{m}}_{\nu}$ or ${\mathfrak{m}}={\mathfrak{m}}_{\rho}$ with $\rho=\rho_{\nu}$. Let ${\mathbb{T}}_{\nu}$ be the localization of $T_{S,K}$ at ${\mathfrak{m}}_{\nu}$, and let
@@
    $\mu$ is sufficiently regular in the sense of [49, 7.18]
@@
The Lan-Suh vanishing conditions are written out more explicitly in [28, §6.10]. We write $H^{q}_{!}({\mathbb{S}},{\mathcal{F}}_{K})$ instead of $H^{q}_{!}(R\Gamma({\mathbb{S}},{\mathcal{F}}_{K}))$.